a number when added to 5, gives the same result as when 2/3 of it is subtracted from 6. what is the number?
Answer by TimothyLamb(4379)

5 + n = 6 - (2/3)n
(3/3)n + (2/3)n = 1
(5/3)n = 1
Answer:
n = 3/5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php